# Not All Data Are Unlearned Equally

### v3

@@ 6 Conclusion
In this work, we test the hypothesis that it should be more difficult to unlearn data that was seen more frequently during pre-training. We provide evidence for this hypothesis across synthetic and real-world setups: 1) We train GPT-2 models from scratch on fake biographies and find that unlearning questions about more frequent biographies is more difficult than other biographies. 2) Extending our results to a real-world scenario, we observe similar trends when unlearning QA pairs from OLMo-7B. Unlearning knowledge that is not frequent in the OLMo pretraining corpus is more easy than frequently encountered data. Additionally, our experiments reveal discrepancies between evaluations of unlearning: While probabilistic evaluations of forget quality and utility display only a minor impact of unlearning on model performance, the impact on performance seems much more negative when using generative evaluations. We hope that our findings inspire the community to develop better evaluation practices and novel, data-dependent LLM unlearning approaches.
@@ Acknowledgments
We would like to thank Fabian David Schmidt for many helpful discussions and feedback on this work. Marius Mosbach is supported by the Mila P2v5 grant and the Mila-Samsung grant. Siva Reddy is supported by the Canada CIFAR AI Chairs program and the NSERC Discovery Grant program.
@@ Appendix A Modified Loss Function for Unlearning
Since our unlearning targets are always QA pairs, and the setup closely resembles Maini et al. (2024) and Fan et al. (2024), we adopt their loss formulation for unlearning. Given a target question-answer pair $(x_{target},y_{target})$ and a retain pair$(x_{retain},y_{retain})$, these papers compute the unlearning loss as:
@@ C.2 Hyperparameters
The following are Questions and Answers. State if the answer to each question is true or false. Question: How many planets are in the solar system? Answer: Eight True or False: True Question: What is the tallest mountain in the world? Answer: Mount Everest True or False: True Question: Where was pizza invented? Answer: France True or False: False Question: Who painted the Mona Lisa? Answer: Leonardo da Vinci True or False: True Question: What color is the gemstone ruby? Answer: Blue True or False: False Question: Are sharks mammals? Answer: No True or False: False Question: Can humans breathe underwater without equipment? Answer: Yes True or False: False Question: Who wrote the novel 1984? Answer: George Orwell True or False: True Question: What is the chemical symbol ’Au’ for? Answer: Aluminum True or False: False Question: What is the primary ingredient in traditional hummus? Answer: Chickpeas True or False: True Question: What is the capital of China? Answer: Beijing True or False:
@@ C.3 OLMo unlearning Plots
Figure 9: Unlearning high-count and low-count capitals with Gradient Ascent on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 10: Unlearning high-count and low-count ZSRE splits with Gradient Ascent on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 11: Unlearning high-count and low-count Books with Gradient Ascent on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) evaluations.
@@
Figure 12: Unlearning high-count and low-count ZSRE splits with SIMNPO on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 13: Unlearning high-count and low-count Books with SIMNPO on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.
@@
Figure 14: Unlearning high-count and low-count capitals with refusal training on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values. Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 15: Unlearning high-count and low-count ZSRE splits with refusal training on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values. Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 16: Unlearning high-count and low-count Books with refusal training on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values. Low frequent instances are forgotten more efficiently across probabilistic (top middle) evaluations.

### v6

@@ 6 Conclusions and Future Work
In this work, we test the hypothesis that it should be more difficult to unlearn data that was seen more frequently during pre-training. We provide evidence for this hypothesis across synthetic and real-world setups: 1) We train GPT-2 models from scratch on fake biographies and find that unlearning questions about more frequent biographies is more difficult than other biographies. 2) Extending our results to a real-world scenario, we observe similar trends when unlearning QA pairs from OLMo-7B. Unlearning knowledge that is not frequent in the OLMo pre-training corpus is more easy than frequently encountered data. Additionally, our experiments reveal discrepancies between evaluations of unlearning: While probabilistic evaluations of forget quality and utility display only a minor impact of unlearning on model performance, the impact on performance seems much more negative when using generative evaluations.
We make the following suggestions to guide future research in LLM unlearning: 1) Unlearning Methodology: As a first step, future work on optimization based unlearning should group data points in the forget set according to their estimated frequency to make sure that a proposed algorithm works well across different samples. To foster research on data-specific unlearning, we need better benchmarks that offer a range of models where training data is known and unlearning splits can be designed carefully. 2) Utility Evaluation: We call for a more comprehensive evaluation of utility following unlearning, given the variation we observed when reporting utility with different benchmarks. We also advocate placing greater emphasis on generative evaluations over likelihood evaluations, since our findings suggest that generative tasks are more sensitive to model degradation.
@@ Acknowledgments
We would like to thank Fabian David Schmidt for many helpful discussions and feedback on this work. We also thank the COLM 2025 reviewers for their constructive feedback. Marius Mosbach is supported by the Mila P2v5 grant and the Mila-Samsung grant. Siva Reddy is supported by the Canada CIFAR AI Chairs program and the NSERC Discovery Grant program.
@@ Appendix A Modified Loss Function for Unlearning
[…]
@@ C.2 Hyperparameters
The following are Questions and Answers. State if the answer to each question is true or false. Question: How many planets are in the solar system? Answer: Eight True or False: True Question: What is the tallest mountain in the world? Answer: Mount Everest True or False: True Question: Where was pizza invented? Answer: France True or False: False Question: Who painted the Mona Lisa? Answer: Leonardo da Vinci True or False: True Question: What color is the gemstone ruby? Answer: Blue True or False: False Question: Are sharks mammals? Answer: No True or False: False Question: Can humans breathe underwater without equipment? Answer: Yes True or False: False Question: Who wrote the novel 1984? Answer: George Orwell True or False: True Question: What is the chemical symbol ’Au’ for? Answer: Aluminum True or False: False Question: What is the primary ingredient in traditional hummus? Answer: Chickpeas True or False: True Question: What is the capital of China?
Answer: Beijing
True or False:
@@ C.3 OLMo unlearning Plots
Appendix D Behind the Scenes
In this section, the authors would like to document the evolution of the paper, all the way from the ”plan of a concept” that AK and MM wrote up to the camera-ready version that the reader sees. Our motivation is to disillusion the linear-progress story that the camera-ready version portrays. The authors thank Benno Krojer for the inspiration.
D.1 Ideation and Experimental Order
The main idea hatched during a Mont-Royal walk: we’d been thinking about the evaluation of unlearning in general, and found two unconvincing arguments in literature: 1) MMLU results were reported before and after unlearning along with the claim that model utility suffered minimal damage: “What are the boundaries of MMLU utility?” 2) Forget set metrics were averaged: “What does 50% unlearning efficacy on a dataset mean?” . Inspired by Peter Hase’s position paper on model beliefs (Hase et al., 2024), we deemed it unlikely that differently strong beliefs can be equally unlearned with one method. When MM was away for EMNLP 2024, AK and MM had a call where OLMo was discussed and the idea of unlearning different buckets of differently exposed data originated. The OLMo part was done first, with a focus on constructing a setup that avoided fine-tuning the model and working on realistic data.
D.2 Stabilizing OLMo unlearning
Initial unlearning experiments using the loss function from Maini et al. (2024) were very unstable: The standard deviation of the plots was so high that one could not derive conclusions when unlearning different target groups. Initially we thought that the instability arose from the choice in retain set, so we invested a lot of effort in ablating it with not much success. Sometime in January 2025, we tried additional regularization with the question part of the retain set, which stabilized the unlearning trends considerably. The idea to add the forget-question to the regularization came up during a discussion with Fabian David Schmidt two weeks before the submission deadline (which meant that all experiments had to be re-run). While the additional term did not change the OLMo results very much, large GPT-2 models showed considerably less utility hits (comparable to the small models) when it was used.
D.3 GPT Unlearning
After we observed differences in unlearning and utility estimation for OLMo, we became interested in studying the effect of scale on our trends. MM was looking at Allen-Zhu & Li (2025) and AK started experimenting with Allen-Zhu & Li (2024). Initially we mimicked the TOFU setup. training models on all BIO and QA instances, then unlearned a subset of the QA samples. This setup was fragile: trends were quite unstable and we hypothesized that verbatim memorization was working against unlearning. After the QA instances were removed from training (and allowed to be learned through generalization), unlearning started to work well and similar trends to OLMo were observed. Scaling and up-sampling also required some attention: we noticed that if we increased the model size beyond 250M, the model had tendencies to memorize the QA training data and all our bigger models did not generalize well. Scaling biographies to create high-count buckets also faced an overfitting problem: Our experiments with 50x, 100x scaling resulted in overfitting to the high-count bucket and poor generalization.
@@ D.3 GPT Unlearning
Figure 10: Unlearning high-count and low-count capitals with Gradient Ascent on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 11: Unlearning high-count and low-count ZSRE splits with Gradient Ascent on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 12: Unlearning high-count and low-count Books with Gradient Ascent on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) evaluations.
@@
Figure 13: Unlearning high-count and low-count ZSRE splits with SIMNPO on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 14: Unlearning high-count and low-count Books with SIMNPO on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values.
@@
Figure 15: Unlearning high-count and low-count capitals with refusal training on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values. Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 16: Unlearning high-count and low-count ZSRE splits with refusal training on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values. Low frequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@
Figure 17: Unlearning high-count and low-count Books with refusal training on OLMo-7B. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values. Low frequent instances are forgotten more efficiently across probabilistic (top middle) evaluations.